# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 2.1. The Jacobson-Cartier condition
Note that the polynomial $f_{p-1}$ is invariant by translation by any element $a\in k$. In fact, we can also write $f(X)=\oplus_{i=0}^{p-1}g_{i}(X)^{p}(X-a)^{i}$, and by comparing the coefficients we find that $f_{p-1}=g_{p-1}$. It is a classical result that $\omega$ is logarithmic if, and only if, $f(X)=f_{p-1}(X)$ [^4] In the case of curves this result is due to Jacobson. The study of the correspondence $f\to f_{p-1}$ as an operation over differential forms over curves is addressed in subsequent work of Tate, that was later generalized to higher dimensions by Cartier. Because of this, such correspondence is often known under the name of Cartier operator. We therefore deem it reasonable to refer to the result on curves as to the “Jacobson-Cartier” condition. We refer to [11, §10, §11] for a discussion of the topic that brings all these different perspectives together.. In our case, that of differential forms over the projective line, this fact has an elementary proof, that we provide below.
@@
Let $\omega=f(X)dX\in\Omega(k(X))$ be such that $f(X)=\frac{1}{P(X)}$ with $P(X)\in k[X]$. Then $\omega=\frac{dF}{F}$ for some $F\in k(X)$ if, and only if, the $p-1$-th derivative $\left({P}^{p-1}\right)^{(p-1)}$ of the polynomial $P^{p-1}$ is equal to $-1$.
@@
By Proposition 2.1, we have that $\omega=\frac{dF}{F}$ if, and only if, $f=f_{p-1}$. Since the $p-1$-th derivative of $f$ satisfies $f^{(p-1)}=(p-1)!f_{p-1}^{p}=-f_{p-1}^{p}$, then it is equivalent to require that $f^{(p-1)}=-f^{p}$. In the case where $f(X)=\frac{1}{P(X)}$, we find $f^{(p-1)}=\left(\frac{P^{p-1}}{P^{p}}\right)^{(p-1)}=\frac{{(P^{p-1})}^{(p-1)}}{P^{p}}.$ On the other hand, $-f^{p}=-\frac{1}{P^{p}}$, and so the condition $f=f_{p-1}$ becomes $\left({P}^{p-1}\right)^{(p-1)}=-1$. ∎
@@ 2.2. The combinatorics of poles of a space $L_{m+1,n}$
The condition that $n$ logarithmic differential forms $\omega_{1},\dots,\omega_{n}$ generate $\Omega$ a space $L_{m+1,n}$ put additional restrictions. Recall that ${\mathcal{P}}(\Omega)$ is the set of poles of at least a differential form in $\Omega$. Pagot proved the following result [8, Lemme 5 and Lemme 6].
@@
The following Corollary is essentially equivalent to [7, Lemme 2.2.2.]. We include here a proof that uses the notation above for the reader’s convenience.
@@
Let $\omega,\omega^{\prime}$ be logarithmic differential forms on ${\mathbb{P}}^{1}_{k}$ with a unique zero at $\infty$ of order $m-1$, and let ${\mathcal{P}}(\omega)$ and ${\mathcal{P}}(\omega^{\prime})$ be the respective set of poles. Then, ${\mathcal{P}}(\omega)={\mathcal{P}}(\omega^{\prime})$ if, and only if $\langle\omega\rangle_{{\mathbb{F}}_{p}}=\langle\omega^{\prime}\rangle_{{\mathbb{F}}_{p}}$.
@@
Suppose by contradiction that there is no $c\in{\mathbb{F}}_{p}^{\times}$ such that $a_{i}=cb_{i}$ for every $i$. Then there is a $j\in{\mathbb{F}}_{p}^{\times}$ such that $a_{i}+jb_{i}=0$ for some but not all $i$’s. Then, the form $\omega+j\omega^{\prime}$ is a non-zero logarithmic differential form with a zero of order at least $m-1$ at infinity, and at the same time it has at most $m$ simple poles. This is not possible, since the degree of any (non-zero) meromorphic differential form on ${\mathbb{P}}^{1}_{k}$ is $-2$. ∎
@@
and in particular it depends only on the class $[\underline{a}]\in{\mathbb{P}}({\mathbb{F}}_{p}^{n})$. To conclude the proof, we count $\sum_{[\underline{a}]\in{\mathbb{P}}({\mathbb{F}}_{p}^{n})}\big{|}{\mathcal{P}}(\omega_{\underline{a}})\cap{\mathcal{P}}(\omega^{\prime})\big{|}$ in two different ways:
@@ 2.3. Frobenius action and étale pullbacks
Let $\omega=\sum_{i=1}^{\lambda p^{n-1}}\frac{a_{i}}{X-x_{i}}dX\in\Omega$ with $a_{i}\in{\mathbb{F}}_{p}^{\times}$. Then $\Phi(\omega)=\sum_{i=1}^{\lambda p^{n-1}}\frac{a_{i}}{X-x_{i}^{p}}dX$ is clearly logarithmic. Moreover, it has a unique zero at $\infty$ because this condition is equivalent to the first line of the equations (2.15). Since ${\mathbb{F}}_{p}$-linearly independent forms are sent to linearly independent forms, we have that $\Phi(\Omega)$ is a space $L_{\lambda p^{n-1},n}$.
@@ 2.4. Known results on spaces $L_{\lambda,1}$
Let $\lambda\in{\mathbb{Z}}$ with $\lambda>1$ and $(\lambda-1,p)=1$. Consider $f(X):=\frac{X^{\lambda-1}-1}{X^{\lambda-1}}$ and the associated logarithmic differential form $\omega:=\frac{df}{f}=(\lambda-1)\frac{dX}{X^{\lambda}-X}$. Then, we have that $\omega=\frac{dX}{P(X)}$ for $P(X)=\frac{X^{\lambda}-X}{\lambda-1}$. There are $\lambda$ simple poles and no zeroes outside $\infty$, then the unique zero at $\infty$ is of order $\lambda-2$ and $\langle\omega\rangle_{{\mathbb{F}_{p}}}$ is a space $L_{\lambda,1}$
It follows from Example 2.18 that spaces $L_{\lambda,1}$ exist for all $(\lambda-1,p)=1$. The converse is also true: if $p|(\lambda-1)$ then by Proposition 2.14 (iii) the leading term of $P^{p-1}$ vanishes, but this implies that $\deg(P)<\lambda$, which is not possible. The paper [4] by Green and Matignon contains more results on spaces $L_{\lambda,1}$, such as a description of all possible spaces $L_{\lambda,1}$ in the case $\lambda<p+1$. A simple but fundamental example that fits in this case is the following:
@@
Let $p\geq 3$ and $f(X):=\prod_{i=1}^{p-1}(X-i1_{k})^{i}\in k[X]$, where $1_{k}\in k$ is the unity of the field $k$. Then, $\Omega:=\langle\frac{df}{f}\rangle_{{\mathbb{F}_{p}}}$ is a space $L_{p-1,1}$. In fact, by construction the non-zero forms in $\Omega$ are logarithmic and their set of (simple) poles is $\{1,2,\dots,p-1\}$. At the pole $i$, the residue of $\omega$ is equal to $i$. We can then verify that the equations (2.15) are satisfied:
@@
Without the assumption $m<p$, a deeper overview of the possible $m+1$-uples of residues $\underline{a}$ (called Hurwitz data) is contained in Henrio’s Ph.D. thesis [5]. It is worth noting that several questions about Hurwitz data remain unanswered (see also [8, §1.1]).
@@
Let $q=p^{t}$ with $t\geq 1$. In [10, Definition 3.2.] a generalization of spaces $L_{m+1,n}$ to the setting of ${\mathbb{F}}_{q}$-vector spaces is introduced. Namely, a space $L_{m+1,n}^{q}$ is defined as a ${\mathbb{F}}_{q}$-vector space of differential forms on ${\mathbb{P}}^{1}_{k}$ whose nonzero elements have simple poles, a unique zero of order $m-1$ at $\infty$ and residues in ${\mathbb{F}}_{q}$. We remark that all the results proved in this section for spaces $L_{m+1,n}$ generalize to spaces $L_{m+1,n}^{q}$. In particular, the key Proposition 2.1 has a natural analogue, and if one introduces the operator $\nabla$ acting on the field $k(X)$ by
@@
as already considered in [10]. The vanishing of this Moore determinant is a necessary and sufficient condition for ${\mathbb{F}}_{q}$-linear dependence of the elements of $\underline{a}$ and key Theorem 4.8 restated in terms of $q$-Moore determinants and spaces $L_{m+1,n}^{q}$ holds with a very similar proof.
However, we have decided not to state our results in this full generality and the case $q=p$ remains the focus of our study: in fact, only in this setting the elements of a space $L_{m+1,n}^{q}$ are logarithmic, while we do not have a similar interpretation otherwise. This property is crucial to solve a concrete question about lifting local actions of elementary abelian groups (see [8, Théorème 11]), while it remains unknown what geometric interpretation or other use might have the existence of a space $L_{m+1,n}^{q}$ for $q$ not prime.
@@ 3.1. Known results on spaces $L_{\lambda p,2}$
We recall the following fundamental result by Pagot ([8, Proposition 7]). A proof is included, which contains elements that are crucial for the main result of this section, as well as for the generalization that we propose in Section 4.
@@
    $\deg(iQ_{1}+jQ_{2})=\lambda$, for every $[i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$
@@
    The $p-1$-th derivative $\left((Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}\right)^{(p-1)}$ of the polynomial $(Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}$ is equal to $-1$.
@@
Suppose that $\Omega$ is a space $L_{\lambda p,2}$. Let us fix a basis $(\omega_{1},\omega_{2})$ of $\Omega$, and remark that the set of poles of an element $i\omega_{1}+j\omega_{2}\in\Omega$ depends only on the corresponding $[i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$. Let us then denote by $X_{[i:j]}$ the set of poles of differential forms in $\Omega$ that are not poles of $i\omega_{1}+j\omega_{2}$. By the results of section 2.2, every $X_{[i:j]}$ consists of $\lambda$ elements and the set consisting of the $X_{[i:j]}$ for all $[i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$ is a partition of the set of $\lambda(p+1)$ poles of $\Omega$. Let us consider the polynomials
@@
$i\omega_{1}+j\omega_{2}=\dfrac{c_{[i:j]}P_{[i:j]}(X)}{P(X)}dX$
for some nonzero constant $c_{[i:j]}\in k$. The condition that $\omega_{1}$ and $\omega_{2}$ generate an ${\mathbb{F}}_{p}$-vector space is then reflected by the condition that
$c_{[i:j]}P_{[i:j]}(X)=ic_{[1:0]}P_{[1:0]}(X)+jc_{[0:1]}P_{[0:1]}(X).$
Since all the $P_{[i:j]}$ are monic polynomials, this means in particular that $c_{[i:j]}=ic_{[1:0]}+jc_{[0:1]}$.
Let us set $a=\frac{c_{[1:0]}}{c_{[0:1]}}$ and note that $a\notin{\mathbb{F}}_{p}$, otherwise we would have $c_{[-1:a]}=0$, which is not possible. As a result, since $k$ is algebraically closed, there is an element $c\in k^{\times}$ satisfying $c^{p}=\frac{1}{c_{[0:1]}\left(a^{p}-a\right)}$. Let us set
@@
Then, $iQ_{2}-jQ_{1}=c(iaP_{[1:0]}+jP_{[0:1]})=c\frac{c_{[i:j]}}{c_{[0:1]}}P_{[i:j]}$, which is a polynomial of degree $\lambda$.
Moreover, up to multiplication by an element of ${\mathbb{F}}_{p}^{\times}$, we have that $\omega_{1}=\frac{Q_{2}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$ and $\omega_{2}=\frac{-Q_{1}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$, as required by condition $(ii)$. Finally, condition $(iii)$ in the statement is also satisfied, by virtue of Corollary 2.2.
Conversely, let us start with $Q_{1}$ and $Q_{2}$ satisfying the three conditions of the proposition and show that they give rise to a space $L_{\lambda p,2}$. First let us show that the differential form $\omega_{1}=\frac{Q_{2}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$ is logarithmic if, and only if $\omega_{2}=\frac{-Q_{1}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$ is logarithmic. We have that
$\big{(}Q_{1}^{p}Q_{2}-Q_{1}Q_{2}^{p}\big{)}^{p}$ $=Q_{1}^{p}Q_{2}\big{(}Q_{1}^{p}Q_{2}-Q_{1}Q_{2}^{p}\big{)}^{p-1}-Q_{1}Q_{2}^{p}\big{(}Q_{1}^{p}Q_{2}-Q_{1}Q_{2}^{p}\big{)}^{p-1}$
$=Q_{1}^{p}Q_{2}^{p}\big{(}Q_{1}^{p}-Q_{1}Q_{2}^{p-1}\big{)}^{p-1}-Q_{1}^{p}Q_{2}^{p}\big{(}Q_{1}^{p-1}Q_{2}-Q_{2}^{p}\big{)}^{p-1}.$
Since the derivative of the left hand side is 0, we have that the $p-1$-th derivative of the right hand side also vanishes, and therefore that $\left(\big{(}Q_{1}^{p}-Q_{1}Q_{2}^{p-1}\big{)}^{p-1}\right)^{(p-1)}=\left(\big{(}Q_{1}^{p-1}Q_{2}-Q_{2}^{p}\big{)}^{p-1}\right)^{(p-1)}$. This implies that we only need to check via Corollary 2.2 that $\omega_{2}$ is logarithmic to ensure that $\omega_{1}$ is logarithmic too. In fact, this stays true if we replace $\omega_{1}$ with $i\omega_{1}+j\omega_{2}$ since $(Q_{1}^{p}Q_{2}-Q_{1}Q_{2}^{p})=Q_{1}^{p}(iQ_{1}+Q_{2})-Q_{1}(iQ_{1}+Q_{2})^{p}$. Hence, it suffices to have condition $(iii)$ for the differential forms $i\omega_{1}+j\omega_{2}$ to be logarithmic for every $i$ and $j$. In particular, these have simple poles. To prove that the ${\mathbb{F}}_{p}$-vector space generated by $\omega_{1}$ and $\omega_{2}$ is of dimension 2, it suffices to remark that, by $(i)$, there are $\lambda$ poles of $\omega_{1}$ that are not poles of $\omega_{2}$, so that $\omega_{2}$ can not be a multiple of $\omega_{1}$. Finally, condition $(ii)$ is enough to ensure that both $\omega_{1}$ and $\omega_{2}$ have $\infty$ as their only zero. ∎

### v2

@@ 2.1. The Jacobson-Cartier condition
Note that the rational function $f_{p-1}$ is invariant by translation by any element $a\in k$. In fact, we can also write $f(X)=\oplus_{i=0}^{p-1}g_{i}(X)^{p}(X-a)^{i}$, and by comparing the coefficients we find that $f_{p-1}=g_{p-1}$. It is a classical result that $\omega$ is logarithmic if, and only if, $f(X)=f_{p-1}(X)$ [^3] In the case of curves this result is due to Jacobson. The study of the correspondence $f\to f_{p-1}$ as an operation over differential forms over curves is addressed in subsequent work of Tate, that was later generalized to higher dimensions by Cartier. Because of this, such correspondence is often known under the name of Cartier operator. We therefore deem it reasonable to refer to the result on curves as to the “Jacobson-Cartier” condition. We refer to [Serre58, §10, §11] for a discussion of the topic that brings all these different perspectives together.. In our case, that of differential forms over the projective line, this fact has an elementary proof, that we provide below.
@@
Let $\omega=f(X)dX\in\Omega(k(X))$ be such that $f(X)=\frac{1}{P(X)}$ with $P(X)\in k[X]$. Then $\omega=\frac{dF}{F}$ for some $F\in k(X)$ if, and only if, the p1p-1-th derivative $\left({P}^{p-1}\right)^{(p-1)}$ of the polynomial $P^{p-1}$ is equal to 1-1.
@@
By Proposition 2.1, we have that $\omega=\frac{dF}{F}$ if, and only if, $f=f_{p-1}$. Since the p1p-1-th derivative of $f$ satisfies $f^{(p-1)}=(p-1)!f_{p-1}^{p}=-f_{p-1}^{p}$, then it is equivalent to require that $f^{(p-1)}=-f^{p}$. In the case where $f(X)=\frac{1}{P(X)}$, we find $f^{(p-1)}=\left(\frac{P^{p-1}}{P^{p}}\right)^{(p-1)}=\frac{{(P^{p-1})}^{(p-1)}}{P^{p}}.$ On the other hand, $f^{p}=-\frac{1}{P^{p}}$, and so the condition $f=f_{p-1}$ becomes $\left({P}^{p-1}\right)^{(p-1)}=-1$. ∎
@@ 2.2. The combinatorics of poles of a space $L_{m+1,n}$
The condition that $n$ logarithmic differential forms $\omega_{1},\dots,\omega_{n}$ generate $\Omega$ a space $L_{m+1,n}$ put combinatorial restrictions on the set ${\mathcal{P}}(\Omega)$ consisting of all the poles of the elements of $\Omega$. In the first part of this section, we rewrite some of the known restrictions using our notation and including alternative proofs for the reader’s convenience. The first is a lemma proved in [Pagot02, Lemme 5 and Lemme 6].
@@
The following Corollary is essentially equivalent to [Pagot02, Lemme 6]. We include here a proof that uses the notation above for the reader’s convenience.
@@
Remark 2.7.
The results above can be interpreted as special cases of a theorem by Yang ([Yang25, Theorem 3.12]), which analyzes the combinatorics of the branch loci of the lifts of $({\mathbb{Z}}/p{\mathbb{Z}})^{n}$-extensions k[[z]]/k[[t]]k[\![z]\!]/k[\![t]\!]. For every $\Omega$ a space $L_{m+1,n}$ with set of poles ${\mathcal{P}}(\Omega)$ we know by [Pagot02, Théorème 11] that there is a lift R[[Z]]/R[[T]]R[\![Z]\!]/R[\![T]\!] of a $({\mathbb{Z}}/p{\mathbb{Z}})^{n}$-extension k[[z]]/k[[t]]k[\![z]\!]/k[\![t]\!] whose branch locus specializes to ${\mathcal{P}}(\Omega)$ and such that the specializations of the branch loci of its $p$-cyclic subextensions are all of the form ${\mathcal{P}}(\omega)$ for $\omega\in\Omega$. Under this correspondence, Yang’s criteria in the equidistant case are equivalent to Lemma 2.3 and Corollary 2.6.
@@
Let $\omega,\omega^{\prime}$ be logarithmic differential forms on ${\mathbb{P}}^{1}_{k}$ with a unique zero at $\infty$ of order m1m-1, and let ${\mathcal{P}}(\omega)$ and ${\mathcal{P}}(\omega^{\prime})$ be the respective set of poles. Then, ${\mathcal{P}}(\omega)={\mathcal{P}}(\omega^{\prime})$ if, and only if $\langle\omega\rangle_{{\mathbb{F}}_{p}}=\langle\omega^{\prime}\rangle_{{\mathbb{F}}_{p}}$.
@@
Suppose by contradiction that there is no $c\in{\mathbb{F}}_{p}^{\times}$ such that $a_{i}=cb_{i}$ for every $i$. Then there is a $j\in{\mathbb{F}}_{p}^{\times}$ such that $a_{i}+jb_{i}=0$ for some but not all $i$’s. Then, the form $\omega+j\omega^{\prime}$ is a non-zero logarithmic differential form with a zero of order at least m1m-1 at infinity, and at the same time it has at most $m$ simple poles. This is not possible, since the degree of any (non-zero) meromorphic differential form on ${\mathbb{P}}^{1}_{k}$ is 2-2. ∎
@@
and in particular it depends only on the class $\underline{a}]\in{\mathbb{P}}({\mathbb{F}}_{p}^{n})$. To conclude the proof, we count $\sum_{[\underline{a}]\in{\mathbb{P}}({\mathbb{F}}_{p}^{n})}\big|{\mathcal{P}}(\omega_{\underline{a}})\cap{\mathcal{P}}(\omega^{\prime})\big|$ in two different ways:
@@ 2.3. Frobenius action and étale pullbacks
Let $\omega=\sum_{i=1}^{\lambda p^{n-1}}\frac{a_{i}}{X-x_{i}}dX\in\Omega$ with $a_{i}\in{\mathbb{F}}_{p}^{\times}$. Then $\Phi(\omega)=\sum_{i=1}^{\lambda p^{n-1}}\frac{a_{i}}{X-x_{i}^{p}}dX$ is clearly logarithmic. Moreover, it has a unique zero at $\infty$ because this condition is equivalent to the one given by equation (2.16). Since ${\mathbb{F}}_{p}$-linearly independent forms are sent to linearly independent forms, we have that $\Phi(\Omega)$ is a space $L_{\lambda p^{n-1},n}$.
@@ 2.4. Known results on spaces $L_{\lambda,1}$
Let $\lambda\in{\mathbb{Z}}$ with $\lambda>1$ and $p\nmid(\lambda-1)$. Consider $f(X):=\frac{X^{\lambda-1}-1}{X^{\lambda-1}}$ and the associated logarithmic differential form $\omega:=\frac{df}{f}=(\lambda-1)\frac{dX}{X^{\lambda}-X}$. Then, we have that $\omega=\frac{dX}{P(X)}$ for $P(X)=\frac{X^{\lambda}-X}{\lambda-1}$. There are $\lambda$ simple poles and no zeroes outside $\infty$, then the unique zero at $\infty$ is of order $\lambda-2$ and $\langle\omega\rangle_{{\mathbb{F}_{p}}}$ is a space $L_{\lambda,1}$
It follows from Example 2.19 that spaces $L_{\lambda,1}$ exist for all $\lambda$ satisfying $p\nmid(\lambda-1)$. The converse is also true: if $p\mid(\lambda-1)$ then by Proposition 2.15 (iii) the leading term of $P^{p-1}$ vanishes, but this implies that $\deg(P)<\lambda$, which is not possible. The paper [GreenMatignon99] contains more results on spaces $L_{\lambda,1}$, such as a description of all possible spaces $L_{\lambda,1}$ in the case $\lambda<p+1$. A simple but fundamental example that fits in this case is the following:
@@
Let $p\geq 3$ and $f(X):=\prod_{i=1}^{p-1}(X-i1_{k})^{i}\in k[X]$, where $1_{k}\in k$ is the unity of the field $k$. Then, $\Omega:=\langle\frac{df}{f}\rangle_{{\mathbb{F}_{p}}}$ is a space $L_{p-1,1}$. In fact, by construction the non-zero forms in $\Omega$ are logarithmic and their set of (simple) poles is $\{1_{k},2\cdot 1_{k},\dots,(p-1)\cdot 1_{k}\}$. At the pole $i\cdot 1_{k}$, the residue of $\omega$ is equal to $i\cdot 1_{k}$. We can then verify that the equations (2.16) are satisfied:
@@
Without the assumption $m<p$, a deeper overview of the possible $m+1$-tuples of residues $\underline{a}$ (called Hurwitz data) is contained in Henrio’s Ph.D. thesis [HenrioThese]. It is worth noting that several questions about Hurwitz data remain unanswered (see also [Pagot02, §1.1]).
@@
Let $q=p^{t}$ with $t\geq 1$. In [FresnelMatignon23, Definition 3.2.] a generalization of spaces $L_{m+1,n}$ to the setting of ${\mathbb{F}}_{q}$-vector spaces is introduced. Namely, a space $L_{m+1,n}^{q}$ is defined as a ${\mathbb{F}}_{q}$-vector space of differential forms on ${\mathbb{P}}^{1}_{k}$ whose nonzero elements have simple poles, a unique zero of order m1m-1 at $\infty$ and residues in ${\mathbb{F}}_{q}$. We remark that all the results proved in this section for spaces $L_{m+1,n}$ generalize to spaces $L_{m+1,n}^{q}$. In particular, the key Proposition 2.1 has a natural analogue, and if one introduces the operator $\nabla$ acting on the field $k(X)$ by
@@
as already considered in [FresnelMatignon23]. The vanishing of this Moore determinant is a necessary and sufficient condition for ${\mathbb{F}}_{q}$-linear dependence of the elements of $\underline{a}$ and key Theorem 4.10 restated in terms of $q$-Moore determinants and spaces $L_{m+1,n}^{q}$ holds with a very similar proof.
However, we have decided not to state our results in this full generality and the case $q=p$ remains the focus of our study: in fact, only in this setting the elements of a space $L_{m+1,n}^{q}$ are logarithmic, while we do not have a similar interpretation otherwise. This property is crucial to solve a concrete question about lifting local actions of elementary abelian groups (see [Pagot02, Théorème 11]), while it remains unknown what geometric interpretation or other use might have the existence of a space $L_{m+1,n}^{q}$ for $q$ not prime.
@@ 3.1. Known results on spaces $L_{\lambda p,2}$
We recall the following fundamental result by Pagot ([Pagot02, Proposition 7]). A proof is included, which contains elements that are crucial for the main result of this section, as well as for the generalization that we propose in Section 4.
@@
    For every $i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$ we have $\deg(iQ_{1}+jQ_{2})=\lambda$
@@
    The p1p-1-th derivative $\left(\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}}\right)^{(p-1)}$ of the polynomial $\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}}$ is equal to 1-1.
@@
Suppose that $\Omega$ is a space $L_{\lambda p,2}$. Let us fix a basis $(\omega_{1},\omega_{2})$ of $\Omega$, and remark that the set of poles of an element $i\omega_{1}+j\omega_{2}\in\Omega$ depends only on the corresponding $i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$. Let us then denote by $X_{[i:j]}$ the set of poles of differential forms in $\Omega$ that are not poles of $i\omega_{1}+j\omega_{2}$. By the results of section 2.2, every $X_{[i:j]}$ consists of $\lambda$ elements and the set consisting of the $X_{[i:j]}$ for all $i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$ is a partition of the set of $\lambda(p+1)$ poles of $\Omega$. Let us consider the polynomials
@@
$i\omega_{1}+j\omega_{2}=\dfrac{c_{ij}P_{[i:j]}(X)}{P(X)}dX$
for nonzero constants $c_{ij}\in k$ that satisfy the condition
$c_{ij}P_{[i:j]}(X)=ic_{10}P_{[1:0]}(X)+jc_{01}P_{[0:1]}(X).$
Since all the $P_{[i:j]}$ are monic polynomials, this means in particular that $c_{ij}=ic_{10}+jc_{01}$. Let us set $a=\frac{c_{10}}{c_{01}}$ and note that $a\notin{\mathbb{F}}_{p}$, otherwise we would have $c_{(p-1)a}=0$, which is not possible, as $\omega_{1}$ and $\omega_{2}$ are assumed ${\mathbb{F}}_{p}$-linearly independent. As a result, since $k$ is algebraically closed, there is an element $c\in k^{\times}$ satisfying $c^{p}=-\frac{1}{c_{01}\left(a^{p}-a\right)}$. Let us set
@@
Then, $iQ_{2}-jQ_{1}=c(iaP_{[1:0]}+jP_{[0:1]})=c\frac{c_{ij}}{c_{01}}P_{[i:j]}$, which is a polynomial of degree $\lambda$.
Moreover, we have $Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}=-c^{p+1}(a^{p}-a)P=\frac{c}{c_{01}}P$ and hence
$\omega_{1}=\frac{Q_{2}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}\;\;\mbox{and}\;\;\omega_{2}=\frac{-Q_{1}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}},$
as required by condition $(ii)$. Among other things, this proves that $\omega_{1}$ is of the form $\frac{1}{P(X)}dX$, and hence it satisfies the hypotheses of Corollary 2.2 (in fact, the multiplication by the element $c\in k^{\times}$ performed earlier on is crucial for having $\omega_{1}$ in such a form). Applying the corollary, we then have that $\left((Q_{1}Q_{2}^{p-1}-Q_{1}^{p})^{p-1}\right)^{(p-1)}=-1$. In order to prove $(iii)$, we only need to show that
(3.2) $\left((Q_{1}Q_{2}^{p-1}-Q_{1}^{p})^{p-1}\right)^{(p-1)}=\left(\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}}\right)^{(p-1)}.$
To do this, we simply note that
$Q_{2}^{p(p-1)}+(Q_{1}Q_{2}^{p-1}-Q_{1}^{p})^{p-1}=\sum_{i=0}^{p}Q_{1}^{i(p-1)}Q_{2}^{(p-i)(p-1)}=\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}},$
and therefore $((Q_{1}Q_{2}^{p-1}-Q_{1}^{p})^{p-1})^{\prime}=\left(\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}}\right)^{\prime}$, from which it follows a fortiori that
$\left((Q_{1}Q_{2}^{p-1}-Q_{1}^{p})^{p-1}\right)^{(p-1)}=\left(\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}}\right)^{(p-1)}=-1.$
Conversely, let us start with $Q_{1}$ and $Q_{2}$ satisfying the three conditions of the proposition and show that they give rise to a space $L_{\lambda p,2}$. By $(ii)$, we write $\omega_{1}=\frac{Q_{2}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$ and $\omega_{2}=\frac{-Q_{1}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$. By $(i)$, there are poles of $\omega_{1}$ that are not poles of $\omega_{2}$, so that $\omega_{2}$ can not be a multiple of $\omega_{1}$ and $\Omega=\langle\omega_{1},\omega_{2}\rangle_{{\mathbb{F}_{p}}}$ has dimension 2. We then note that any non-zero element $\omega\in\Omega$ can be written as $\omega=\frac{R_{2}dX}{R_{1}R_{2}^{p}-R_{1}^{p}R_{2}}$ for some pair of polynomials $(R_{1},R_{2})=(Q_{1},Q_{2})M$ obtained multiplying $(Q_{1},Q_{2})$ by a suitable matrix $M\in GL_{2}({\mathbb{F}}_{p})$. It follows that $\omega$ has a unique zero at $\infty$. We leave the reader with the task to check that $\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}}=\frac{R_{2}^{p^{2}-1}-R_{1}^{p^{2}-1}}{R_{2}^{p-1}-R_{1}^{p-1}}$. Combining this with condition $(iii)$ we get that $\left(\frac{R_{2}^{p^{2}-1}-R_{1}^{p^{2}-1}}{R_{2}^{p-1}-R_{1}^{p-1}}\right)^{(p-1)}=-1$. By applying equation (3.2) we find that $\left((R_{1}R_{2}^{p-1}-R_{1}^{p})^{p-1}\right)^{(p-1)}=-1$, which by Corollary 2.2 implies that $\omega$ is logarithmic for every $\omega\in\Omega$. ∎